Circular No. 7133

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1999bc AND 1999bd P. Nugent, Lawrence Berkeley Laboratory, on behalf of the Supernova

Cosmology Project and the Near-Earth Asteroid Tracking project (cf. IAUC 7125, 7128), reports two supernovae discoveries on unfiltered CCD frames:

SN 1999 UT α_{2000} δ_{2000} Mag. Offset
1999bc Feb. 19 8^{h}28^{m}31.^{s}08 +17^{o}28′15′′.4 18.9: 4′′.2 W, 17′′.1 N
1999bd Feb. 19 9 30 29.17 +16 26 07.8 17.7:  ?

A CCD spectrum (range 420–850 nm) of SN 1999bc, located in UGC 4433, taken by G. Aldering on Mar. 5 UT at the MDM 2.4-m telescope, showed a few narrow emission features (463, 537, and 586 nm) ∼ 4000 km/s across, along with some contamination from a superposed H II region. A CCD spectrum (range 365–930 nm) taken by M. M. Phillips on Mar. 22 with the Las Campanas 2.5-m duPont telescope shows that these features have broadened somewhat over this time period. Phillips classifies SN 1999bc as a type-Ic supernova at a redshift of 0.022, the spectrum being similar to that of SN 1983V at ∼ 30 days after peak brightness.

SN 1999bd has no apparent host on the discovery images (limiting mag ∼ 19). A CCD spectrum obtained at the ARC 3.5-m telescope on Mar. 10 by G. Richards, H. Newberg, and S. Kent showed a blue continuum with several narrow features and Hα emission, strongly absorbed by the atmospheric A band. Phillips took another spectrum of this object (range 365–930) on Mar. 22, and he classifies it as a type-II supernova at a redshift of 0.151. Once again, several narrow features are present in the spectrum, along with a weak, broad component visible below Hα.

XTE J1723–376 F. E. Marshall, Goddard Space Flight Center (GSFC); Y. Ueda, Insti-

tute of Space and Astronautical Science; and C. B. Markwardt, National Academy of Sciences/National Research Council and GSFC, report an improved position for XTE J1723–376 (IAUC 7103): α = 17^{h}23^{m}38^{s}, δ = −37^{o}39′.7 (equinox 2000.0; estimated uncertainty 0′.5, from an ASCA observation on Mar. 4 UT). The x-ray spectrum is typical of low-mass x-ray binaries, and the absorption is ∼ 4 × 10^{22} cm^{−2}. Type-1 x-ray bursts were observed with the RXTE PCA on Feb. 3.942 and 4.055, with peak intensities of 0.51 and 0.43 Crab. These results indicate that the source is at a distance comparable to that of the Galactic Center. The most recent RXTE observation shows a persistent flux of 0.05 Crab on Mar. 22.9.

1999 March 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green